OpenStudy (jordanboss):

help me with this eqaution it is 8=1/3(3.14)X(16)h

OpenStudy (wholock):

Hello! Now, I'm going to need a little clarification, the equation is this, right: \[8 = \frac{ 1 }{ 3 }(3.14 \times 16)h\]

OpenStudy (jordanboss):

yes

OpenStudy (wholock):

Okay, well first, handle the parentheses. 3.14 * 16 = 50.24 Now, distribution property, 50.24 * (1/3) = 16.7 (roughly) Now you have 8 = 16.7h Divide both sides by 16.7 to get: h =(roughly) .48

OpenStudy (jordanboss):

thank you

OpenStudy (wholock):

Sure thing.